# Communication-Efficient Federated Group Distributionally Robust Optimization

### v1

@@ Abstract
Federated learning faces challenges due to the heterogeneity in data volumes and distributions at different clients, which can compromise model generalization ability to various distributions. Existing approaches to address this issue based on group distributionally robust optimization (GDRO) often lead to high communication and sample complexity. To this end, this work introduces algorithms tailored for communication-efficient Federated Group Distributionally Robust Optimization (FGDRO). Our contributions are threefold: Firstly, we introduce the FGDRO-CVaR algorithm, which optimizes the average top-K losses while reducing communication complexity to $O(1/\epsilon^{4})$, where $\epsilon$ denotes the desired precision level. Secondly, our FGDRO-KL algorithm is crafted to optimize KL regularized FGDRO, cutting communication complexity to $O(1/\epsilon^{3})$. Lastly, we propose FGDRO-KL-Adam to utilize Adam-type local updates in FGDRO-KL, which not only maintains a communication cost of $O(1/\epsilon^{3})$ but also shows potential to surpass SGD-type local steps in practical applications. The effectiveness of our algorithms has been demonstrated on a variety of real-world tasks, including natural language processing and computer vision.
@@ 4 FGDRO-CVaR
The formulation (2) replaces the constrained high-dimensional vector $\mathbf{p}$ with a single unconstrained scalar variable $s$. However, this adjustment introduces new challenges due to the compositional structure and the non-smooth nature of the outer function $f$. As a result, it is biased to estimate subgradient $\partial f(g_{i}(\mathbf{w})-s)\nabla g_{i}(\mathbf{w})$ using a batch of samples. To address this, it is common practice to create an accurate estimate of $g_{i}(\mathbf{w})$ [24, 22, 64, 65, 32]. Specifically, we employ a moving average $\mathbf{u}$ as our accurate estimate:
@@ 7 Experiments
We use Pile [15], CivilComments [5], Camelyon17 [1], iWildCam2020 [3], and Poverty [72]. For Pile, we preprocess it as [70], for the others, we use the preprocessed version by [37]. We utilized natural data splits where data from the same hospital, web source, location, or demographic group were placed on the same client machine. These experiments have involved with highly imbalanced number of data on clients. Data statistics are presented in the Appendix E. Additiona experiments on Cifar 10 with Dirichlet distributions over 100 clients are reported in Appendix G.
@@ B.2 Proof of Lemma 32
$\begin{split}&\mathbb{E}\|\mathbf{m}^{r}_{i,\tau}-\bar{\mathbf{m}}^{r}_{\tau}\|^{2}=\mathbb{E}\bigg{\|}\left(\sum\limits_{\tau=1}^{t}(1-\beta_{3})^{\tau-1}\beta_{3}\frac{1}{v^{r}_{i,\tau}}g(u^{r}_{i,\tau})\nabla\ell(\mathbf{w}^{r}_{i,\tau-1};\mathbf{z}^{r}_{\tau,t})+(1-\beta_{3})^{\tau}\bar{\mathbf{m}}^{r}_{0}\right)\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\left(\sum\limits_{\tau=1}^{t}(1-\beta_{3})^{\tau-1}\beta_{3}\frac{1}{N}\sum\limits_{k=1}^{N}\frac{1}{v^{r}_{k,t}}g(u^{r}_{k,t})\nabla\ell(\mathbf{w}^{r}_{k,t-1};\mathbf{z}^{r}_{k,t})+(1-\beta_{3})^{\tau}\bar{\mathbf{m}}^{r}_{0}\right)\bigg{\|}^{2}\\ &\leq 2\mathbb{E}\|\sum\limits_{\tau=1}^{t}(1-\beta_{3})^{\tau-1}\beta_{3}\frac{1}{v^{r}_{i,\tau}}g(u^{r}_{i,\tau})\nabla\ell(\mathbf{w}^{r}_{i,\tau-1};\mathbf{z}^{r}_{\tau,t})\|^{2}\\ &~{}~{}~{}+2\mathbb{E}\|\sum\limits_{\tau=1}^{t}(1-\beta_{3})^{\tau-1}\beta_{3}\frac{1}{N}\sum\limits_{k=1}^{N}\frac{1}{v^{r}_{k,t}}g(u^{r}_{k,t})\nabla\ell(\mathbf{w}^{r}_{k,t-1};\mathbf{z}^{r}_{k,t})\|^{2}\\ &\leq 2t\sum\limits_{\tau=1}^{t}\beta_{3}^{2}\mathbb{E}[\|\frac{1}{v^{r}_{i,\tau}}\|^{2}\|g(u^{r}_{i,\tau})\|^{2}\|\nabla\ell(\mathbf{w}^{r}_{i,\tau-1};\mathbf{z}^{r}_{i,\tau})\|^{2}]\\ &+2t\sum\limits_{\tau=1}^{t}\beta_{3}^{2}\frac{1}{N}\sum\limits_{k=1}^{N}\mathbb{E}[\|\frac{1}{v^{r}_{k,t}}\|^{2}\|g(u^{r}_{k,\tau})\|^{2}\|\nabla\ell(\mathbf{w}^{r}_{k,t-1};\mathbf{z}^{r}_{i,\tau})\|^{2}]\\ &\leq 8I^{2}\beta_{3}^{2}C_{1}^{2}\bigg{(}\mathbb{E}\|\nabla\ell(\mathbf{w}^{r}_{i,\tau-1};\mathbf{z}^{r}_{i,\tau})-\nabla\ell(\mathbf{w}^{r}_{i,\tau-1};\mathcal{D}_{i})\|^{2}+\mathbb{E}\|\nabla\ell(\mathbf{w}^{r}_{i,\tau-1};\mathcal{D}_{i})\|^{2}\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{1}{N}\sum\limits_{k=1}^{N}(\mathbb{E}\|\nabla\ell(\mathbf{w}^{r}_{k,\tau-1};\mathbf{z}^{r}_{k,\tau})-\nabla\ell(\mathbf{w}^{r}_{k,\tau-1};\mathcal{D}_{k})\|^{2}+\mathbb{E}\|\nabla\ell(\mathbf{w}^{r}_{k,\tau-1};\mathcal{D}_{k})\|^{2})\bigg{)}\\ &\leq 16I^{2}\beta_{3}^{2}C_{1}^{2}(\sigma^{2}+C_{\ell}^{2}),\end{split}$ (36)
@@ B.3 Proof of Lemma 33
$\begin{split}&(A)\leq\left(1+\frac{\beta_{3}}{2}\right)\bigg{\|}(1-\beta_{3})(\bar{\mathbf{m}}^{r}_{t-1}-\nabla F(\bar{\mathbf{w}}^{r}_{t-1}))\\ &~{}~{}~{}~{}~{}~{}+\beta_{3}\left(\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(u^{r}_{i,t-1}/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathcal{D}_{i})-\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(\ell(\bar{\mathbf{w}}^{r}_{t-1};\mathcal{D}_{i})/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathcal{D}_{i})\right)\\ &~{}~{}~{}~{}~{}~{}+\beta_{3}\left(\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(\ell(\bar{\mathbf{w}}^{r}_{t-1};\mathcal{D}_{i})/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathcal{D}_{i})-\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{\bar{v}^{r}_{t}}\exp(\ell(\bar{\mathbf{w}}^{r}_{t-1};\mathcal{D}_{i})/\lambda)\nabla\ell(\bar{\mathbf{w}}^{r}_{t-1};\mathcal{D}_{i})\right)\\ &~{}~{}~{}~{}~{}~{}+\beta_{3}\left(\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{\bar{v}^{r}_{t}}\exp(\ell(\bar{\mathbf{w}}^{r}_{t-1};\mathcal{D}_{i})/\lambda)\nabla\ell(\bar{\mathbf{w}}^{r}_{t-1};\mathcal{D}_{i})-\nabla F(\bar{\mathbf{w}}^{r}_{t-1})\right)\bigg{\|}^{2},\\ &~{}~{}~{}+\left(1+\frac{\beta_{3}}{2}\right)\left\|\beta_{3}\left(\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(u^{r}_{i,t-1}/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathbf{z}^{r}_{i,t})-\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(u^{r}_{i,t-1}/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathcal{D}_{i})\right)\right\|^{2}\\ &~{}~{}~{}+(1+\frac{2}{\beta_{3}})\beta_{3}\left\|\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(u^{r}_{i,t}/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathbf{z}^{r}_{i,t})-\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(u^{r}_{i,t-1}/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathbf{z}^{r}_{i,t})\right\|^{2},\end{split}$ (41)
which leads to
@@
Then it follows that
@@ Appendix E Statistics of Datasets
Appendix F Running Time
Running time is reported in Tabel 6. Each algorithm was run on a high performance cluster where each machine uses a NVIDIA A100 GPU.
Pile CivilComments Camelyon17 iWildCam2020 PovertyMap
FedAvg 29.77 3.05 4.59 12.90 1.77
FedAdam 35.39 3.13 6.01 12.91 3.22
DRFA 34.12 3.27 4.94 12.95 2.25
DR-DSGD 25.26 4.01 6.26 13.02 4.71
FGDRO-CVAR 34.07 4.03 7.19 13.01 5.02
FGDRO-KL 34.92 3.65 7.23 14.57 5.42
FGDRO-KL-Adam 35.98 3.80 7.54 14.82 5.50
Appendix G Experiments on Cifar 10
@@ 38. Answer: [Yes]
  39. Justification: Computing resource and time of execution are reported in Appendix F.

### v2

@@ Abstract
Federated learning faces challenges due to the heterogeneity in data volumes and distributions at different clients, which can compromise model generalization ability to various distributions. Existing approaches to address this issue based on group distributionally robust optimization (GDRO) often lead to high communication and sample complexity. To this end, this work introduces algorithms tailored for communication-efficient Federated Group Distributionally Robust Optimization (FGDRO). Our contributions are threefold: Firstly, we introduce the FGDRO-CVaR algorithm, which optimizes the average top-K losses while reducing communication complexity to $O(1/\epsilon^{4})$, where $\epsilon$ denotes the desired precision level. Secondly, our FGDRO-KL algorithm is crafted to optimize KL regularized FGDRO, cutting communication complexity to $O(1/\epsilon^{3})$. Lastly, we propose FGDRO-KL-Adam to to utilize Adam-type local updates in FGDRO-KL, which not only maintains a communication cost of $O(1/\epsilon^{3})$ but also shows potential to surpass SGD-type local steps in practical applications. The effectiveness of our algorithms has been demonstrated on a variety of real-world tasks, including natural language processing and computer vision.
@@ 4 FGDRO-CVaR
The formulation (2) replaces the constrained high-dimensional vector $\mathbf{p}$ with a single unconstrained scalar variable $s$. However, this adjustment introduces new challenges due to the compositional structure and the non-smooth nature of the outer function $f$. As a result, it is biased to estimate subgradient $\partial f(g_{i}(\mathbf{w})-s)\nabla g_{i}(\mathbf{w})$ using a batch of samples. To address this, it is common practice to create an accurate estimate of $g_{i}(\mathbf{w})$ [67, 24, 22, 66, 32, 55, 57]. Specifically, we employ a moving average $\mathbf{u}$ as our accurate estimate:
@@ 7 Experiments
We use Pile [15], CivilComments [5], Camelyon17 [1], iWildCam2020 [3], and Poverty [74]. For Pile, we preprocess it as [72], for the others, we use the preprocessed version by [37]. We utilized natural data splits where data from the same hospital, web source, location, or demographic group were placed on the same client machine. These experiments have involved with highly imbalanced number of data on clients. Data statistics are presented in the Appendix E. Additiona experiments on Cifar 10 with Dirichlet distributions over 100 clients are reported in Appendix F.
@@ 9 Impact Statements
Acknowledgments and Disclosure of Funding
We appreciate the feedback provided by the anonymous reviewers. This work was partially supported by the National Science Foundation Career Award 2246753, the National Science Foundation Award 2246757, 2246756 and 2306572.
@@ B.2 Proof of Lemma 32
$\begin{split}&\mathbb{E}\|\mathbf{m}^{r}_{i,\tau}-\bar{\mathbf{m}}^{r}_{\tau}\|^{2}=\mathbb{E}\bigg{\|}\left(\sum\limits_{\tau=1}^{t}(1-\beta_{3})^{\tau-1}\beta_{3}\frac{1}{v^{r}_{i,\tau}}g(u^{r}_{i,\tau})\nabla\ell(\mathbf{w}^{r}_{i,\tau-1};\mathbf{z}^{r}_{\tau,t})+(1-\beta_{3})^{\tau}\bar{\mathbf{m}}^{r}_{0}\right)\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\left(\sum\limits_{\tau=1}^{t}(1-\beta_{3})^{\tau-1}\beta_{3}\frac{1}{N}\sum\limits_{k=1}^{N}\frac{1}{v^{r}_{k,t}}g(u^{r}_{k,t})\nabla\ell(\mathbf{w}^{r}_{k,t-1};\mathbf{z}^{r}_{k,t})+(1-\beta_{3})^{\tau}\bar{\mathbf{m}}^{r}_{0}\right)\bigg{\|}^{2}\\ &\leq 2\mathbb{E}\|\sum\limits_{\tau=1}^{t}(1-\beta_{3})^{\tau-1}\beta_{3}\frac{1}{v^{r}_{i,\tau}}g(u^{r}_{i,\tau})\nabla\ell(\mathbf{w}^{r}_{i,\tau-1};\mathbf{z}^{r}_{\tau,t})\|^{2}\\ &~{}~{}~{}+2\mathbb{E}\|\sum\limits_{\tau=1}^{t}(1-\beta_{3})^{\tau-1}\beta_{3}\frac{1}{N}\sum\limits_{k=1}^{N}\frac{1}{v^{r}_{k,t}}g(u^{r}_{k,t})\nabla\ell(\mathbf{w}^{r}_{k,t-1};\mathbf{z}^{r}_{k,t})\|^{2},\end{split}$ (36)
which yields
$\begin{split}&\mathbb{E}\|\mathbf{m}^{r}_{i,\tau}-\bar{\mathbf{m}}^{r}_{\tau}\|^{2}\\ &\leq 2t\sum\limits_{\tau=1}^{t}\beta_{3}^{2}\mathbb{E}[\|\frac{1}{v^{r}_{i,\tau}}\|^{2}\|g(u^{r}_{i,\tau})\|^{2}\|\nabla\ell(\mathbf{w}^{r}_{i,\tau-1};\mathbf{z}^{r}_{i,\tau})\|^{2}]\\ &+2t\sum\limits_{\tau=1}^{t}\beta_{3}^{2}\frac{1}{N}\sum\limits_{k=1}^{N}\mathbb{E}[\|\frac{1}{v^{r}_{k,t}}\|^{2}\|g(u^{r}_{k,\tau})\|^{2}\|\nabla\ell(\mathbf{w}^{r}_{k,t-1};\mathbf{z}^{r}_{i,\tau})\|^{2}]\\ &\leq 8I^{2}\beta_{3}^{2}C_{1}^{2}\bigg{(}\mathbb{E}\|\nabla\ell(\mathbf{w}^{r}_{i,\tau-1};\mathbf{z}^{r}_{i,\tau})-\nabla\ell(\mathbf{w}^{r}_{i,\tau-1};\mathcal{D}_{i})\|^{2}+\mathbb{E}\|\nabla\ell(\mathbf{w}^{r}_{i,\tau-1};\mathcal{D}_{i})\|^{2}\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{1}{N}\sum\limits_{k=1}^{N}(\mathbb{E}\|\nabla\ell(\mathbf{w}^{r}_{k,\tau-1};\mathbf{z}^{r}_{k,\tau})-\nabla\ell(\mathbf{w}^{r}_{k,\tau-1};\mathcal{D}_{k})\|^{2}+\mathbb{E}\|\nabla\ell(\mathbf{w}^{r}_{k,\tau-1};\mathcal{D}_{k})\|^{2})\bigg{)}\\ &\leq 16I^{2}\beta_{3}^{2}C_{1}^{2}(\sigma^{2}+C_{\ell}^{2}),\end{split}$ (37)
@@ B.3 Proof of Lemma 33
$\begin{split}&(A)\leq\left(1+\frac{\beta_{3}}{2}\right)\bigg{\|}(1-\beta_{3})(\bar{\mathbf{m}}^{r}_{t-1}-\nabla F(\bar{\mathbf{w}}^{r}_{t-1}))\\ &~{}~{}~{}~{}~{}~{}+\beta_{3}\left(\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(u^{r}_{i,t-1}/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathcal{D}_{i})-\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(\ell(\bar{\mathbf{w}}^{r}_{t-1};\mathcal{D}_{i})/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathcal{D}_{i})\right)\\ &~{}~{}~{}~{}~{}~{}+\beta_{3}\left(\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(\ell(\bar{\mathbf{w}}^{r}_{t-1};\mathcal{D}_{i})/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathcal{D}_{i})-\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{\bar{v}^{r}_{t}}\exp(\ell(\bar{\mathbf{w}}^{r}_{t-1};\mathcal{D}_{i})/\lambda)\nabla\ell(\bar{\mathbf{w}}^{r}_{t-1};\mathcal{D}_{i})\right)\\ &~{}~{}~{}~{}~{}~{}+\beta_{3}\left(\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{\bar{v}^{r}_{t}}\exp(\ell(\bar{\mathbf{w}}^{r}_{t-1};\mathcal{D}_{i})/\lambda)\nabla\ell(\bar{\mathbf{w}}^{r}_{t-1};\mathcal{D}_{i})-\nabla F(\bar{\mathbf{w}}^{r}_{t-1})\right)\bigg{\|}^{2},\\ &~{}~{}~{}+\left(1+\frac{\beta_{3}}{2}\right)\left\|\beta_{3}\left(\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(u^{r}_{i,t-1}/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathbf{z}^{r}_{i,t})-\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(u^{r}_{i,t-1}/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathcal{D}_{i})\right)\right\|^{2}\\ &~{}~{}~{}+(1+\frac{2}{\beta_{3}})\beta_{3}\left\|\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(u^{r}_{i,t}/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathbf{z}^{r}_{i,t})-\frac{1}{N}\sum\limits_{i=1}^{N}\frac{1}{v^{r}_{i,t}}\exp(u^{r}_{i,t-1}/\lambda)\nabla\ell(\mathbf{w}^{r}_{i,t-1};\mathbf{z}^{r}_{i,t})\right\|^{2}.\end{split}$ (42)
Then it leads to
@@
Thus, it follows that
@@ Appendix E Statistics of Datasets
Appendix F Experiments on Cifar 10
@@
Appendix G Running Time
Running time is reported in Tabel 7. Each algorithm was run on a high performance cluster where each machine uses a NVIDIA A100 GPU.
Pile CivilComments Camelyon17 iWildCam2020 PovertyMap
FedAvg 29.77 3.05 4.59 12.90 1.77
FedAdam 35.39 3.13 6.01 12.91 3.22
DRFA 34.12 3.27 4.94 12.95 2.25
DR-DSGD 25.26 4.01 6.26 13.02 4.71
FGDRO-CVAR 34.07 4.03 7.19 13.01 5.02
FGDRO-KL 34.92 3.65 7.23 14.57 5.42
FGDRO-KL-Adam 35.98 3.80 7.54 14.82 5.50
@@ 38. Answer: [Yes]
  39. Justification: Computing resource and time of execution are reported in Appendix G.